Oxygen (O2) gas is observed to diffuse through a capillary for 18 seconds. How long will it take for methane gas (CH4) to diffuse through the same orifice under identical conditions of temperature and pressure?


1
Expert's answer
2022-01-20T16:06:03-0500

It is based on Graham's law of diffusion rates: "The rate of diffusion of a gas is inversely proportional to the square root of it's molar mass."

The formula for comparing the diffusion rates of two gases:

r1/r2 = (M2/M1)1/2

​r stands for rate of diffusion and M stands for molar mass of the gas and the subscripts 1 and 2 are referring two different gases.

Let's say 1 is for oxygen and 2 is for methane. Molar mass of oxygen is 32 and molar mass of methane is 16.

Rate is how much volume diffused per unit time, so:

(V1/t1) / (V2/t2) = (M2/M1)1/2

Since the volumes are same, the formula could be written as:

t2/t1 = (M2/M1)1/2

t2/18 sec = (16/32)1/2

t2 = 12.7 sec